# Better Instruction-Following Through Minimum Bayes Risk

### v3

@@ 6 Conclusion
In an effort to make our work reproducible, we document all prompts (Appendices E and B), as well as training and inference hyperparameters (Appendix I) used throughout our experiments. We also include version information for all API-based LLMs (Appendix B), and choose to use open-source models (the Llama2, Llama3, Prometheus-2 and JudgeLM families) where possible.
@@ A.1 Additional MBR Utility Metrics
Llama2-7B Llama2-70B Avg. $\Delta$
@@
We experiment with two additional ROUGE variants and two additional dense embedders as utility metrics. The two ROUGE variants are ROUGE-2 and ROUGE-L, which detect bigram overlap and longest co-occurring n-gram overlap respectively. The two dense embedders are NVEmbed-v2 (Lee et al., 2024) and Nomic-Text-v1.5 (Nussbaum et al., 2024), which, like SFR Embedder, are strong, long-context dense embedders that rank highly on the MTEB leaderboard Muennighoff et al. (2023).
@@ A.2 Comparison with Universal Self-Consistency
The limited context lengths of LLMs poses a significant challenge when using USC, as it requires fitting all $N_{\mathrm{cand}}=30$ samples into a single prompt. In contrast, MBR decoding only requires fitting two outputs into a single prompt as utility is computed pairwise. As our existing choice of models have limited context lengths (4096 tokens for Llama2, 8192 tokens for Llama3) and our outputs can be long (up to 1024 tokens), we are unable to assess USC on equal footing with MBR decoding using these models without significantly reducing $N_{\mathrm{cand}}$. In order to facilitate a fair comparison, we therefore use the Llama-3.1 models (Dubey et al., 2024) in place of the Llama2 and Llama3 models for this experiment. The Llama-3.1 models possess context lengths of 128k, thereby allowing us to fit all $N_{\mathrm{cand}}$ samples into a single to prompt as required. Our USC prompt is as follows:
@@
Llama3.1-8B Llama3.1-70B Avg. $\Delta$
@@
Table 6: AlpacaEval 2.0 win rates (%) for Llama3.1 models with greedy decoding, USC and MBR decoding with Prometheus, along with the average win rate differences compared to greedy decoding across all models (denoted Avg. $\Delta$).
@@ A.3 Sampling Baselines
We conduct additional baseline experiments with top-$p$ (Holtzman et al., 2020) and top-$k$ (Fan et al., 2018) sampling. We use these sampling methods to directly obtain a final output, and do not explore using replacing temperature sampling with these sampling strategies for generating the hypothesis set - we leave this to future work.
Llama2-7B Llama2-70B Avg. $\Delta$
@@
We find that top-$p$ and top-$k$ sampling improves over greedy decoding, and can achieve performance close to that of MBR decoding with ROUGE. The improvements are nonetheless much smaller than MBR decoding with Prometheus. Nonetheless, these results demonstrate that top-$p$ and top-$k$ sampling could be used to produce hypothesis sets containing higher quality candidates, which could in turn improve downstream MBR decoding performance. We believe this is an exciting avenue for future work.
@@ Appendix D AlpacaEval Categories
Appendix E Llama3 as an MBR and BoN Utility Metric
We experiment with using Llama3-70b-Instruct as an MBR and BoN utility metric in Section 3.3.1. This entails prompting the model to act as either a reference-based or reference-free evaluator.
Our prompt for single-turn reference-based evaluation is as follows:
@@ G.3 MBR Self-Training with ROUGE
G.4 Alternative Pair Selection Strategies
@@ G.5 Performance by Category
We replicate the question category analysis described in Section 3.3 and Appendix D for our Prometheus MBR self-trained models and report results in Figure 4. We find that performance relative to the sft models improves across almost all question categories, with performance on writing based categories improving more than on reasoning based categories. Performance on mathematical reasoning remains unchanged for the 7B model and decreases for the 13B model.
@@ Appendix H Human Study
Table 14: Head-to-head evaluation of Prometheus MBR and Prometheus BoN vs. greedy decoding for Llama2-70b conducted on the AlpacaEval dataset by human evaluators.
@@
We conducted four head-to-head experiments. The first two experiments, corresponding to experiments in Section 3, were between Prometheus MBR vs. greedy decoding and Prometheus BoN vs. greedy decoding with Llama2-70b. The next two experiments, corresponding to experiments in Section 4, were between 13B dpo-3-MBR vs. sft and dpo-3-BoN vs. sft. We used 60 samples for each experiment. Volunteers were given an even distribution of samples from each experiment.
@@
Appendix I Training and Inference Hyperparameters
@@
Our SFT and DPO hyperparameters for our self-training experiments in Section 4.2 are provided in Tables 16 and 17. We use bf16 mixed precision training with 8xA100 GPUs for all experiments.

### v4

@@ 6 Conclusion
In an effort to make our work reproducible, we document all prompts (Appendices E and B), as well as training and inference hyperparameters (Appendix K) used throughout our experiments. We also include version information for all API-based LLMs (Appendix B), and choose to use open-source models (the Llama-2, Llama-3, Prometheus-2 and JudgeLM families) where possible.
Acknowledgements
AB was supported by a grant from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739. PF was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI) and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ A.1 Additional MBR Utility Metrics
Llama-2-7B Llama-2-70B Avg. $\Delta$
@@
We experiment with two additional ROUGE variants and two additional dense embedders as utility metrics. The two ROUGE variants are ROUGE-2 and ROUGE-L, which detect bigram overlap and longest co-occurring n-gram overlap respectively. The two dense embedders are NVEmbed-v2 (Lee et al., 2024) and Nomic-Text-v1.5 (Nussbaum et al., 2024), which, like SFR Embedder, are strong, long-context dense embedders that rank highly on the MTEB leaderboard (Muennighoff et al., 2023).
@@ A.2 Comparison with Universal Self-Consistency
The limited context lengths of LLMs poses a significant challenge when using USC, as it requires fitting all $N_{\mathrm{cand}}=30$ samples into a single prompt. In contrast, MBR decoding only requires fitting two outputs into a single prompt as utility is computed pairwise. As our existing choice of models have limited context lengths (4096 tokens for Llama-2, 8192 tokens for Llama-3) and our outputs can be long (up to 1024 tokens), we are unable to assess USC on equal footing with MBR decoding using these models without significantly reducing $N_{\mathrm{cand}}$. In order to facilitate a fair comparison, we therefore use the Llama-3.1 models (Dubey et al., 2024) in place of the Llama-2 and Llama-3 models for this experiment. The Llama-3.1 models possess context lengths of 128k, thereby allowing us to fit all $N_{\mathrm{cand}}$ samples into a single to prompt as required. Our USC prompt is as follows:
@@
Llama-3.1-8B Llama-3.1-70B Avg. $\Delta$
@@
Table 6: AlpacaEval 2.0 win rates (%) for Llama-3.1 models with greedy decoding, USC and MBR decoding with Prometheus, along with the average win rate differences compared to greedy decoding across all models (denoted Avg. $\Delta$).
@@ A.3 Sampling Baselines
We conduct additional baseline experiments with top-$p$ (Holtzman et al., 2020) and top-$k$ (Fan et al., 2018) sampling. We use these sampling methods to directly obtain a final output, and do not explore replacing standard temperature sampling with these sampling strategies for generating the hypothesis set - we leave this to future work.
Llama-2-7B Llama-2-70B Avg. $\Delta$
@@
We find that top-$p$ and top-$k$ sampling improves over greedy decoding, and can achieve performance close to that of MBR decoding with ROUGE. The improvements are nonetheless much smaller than MBR decoding with Prometheus. Nonetheless, these results demonstrate that top-$p$ and top-$k$ sampling could be used to produce hypothesis sets containing higher quality candidates, which could in turn improve downstream MBR decoding performance.
@@ A.4 Generation Lengths
A.5 Linguistic Features
Bullet List Numbered List Both Neither
Greedy 14.2 29.1 26.4 30.2
Prometheus BoN 13.6 29.8 26.0 30.8
Prometheus MBR 13.9 30.1 28.1 27.8
Table 8: Percentage of responses to AlpacaEval containing bullet lists, numbered lists, both bullet and numbered lists, and neither bullet nor numbered lists. These responses were generated by both Llama-3-8b-Instruct and Llama-3-70b-Instruct using various decoding strategies. We find that there is little difference in the formatting of MBR-generated, BoN-generated and greedy-decoded outputs.
TTR FK
Greedy 0.514 12.13
Prometheus BoN 0.520 12.40
Prometheus MBR 0.521 12.24
Table 9: Averaged type token ratios (TTR) and Flesch Kincaid readability scores (FK) across outputs generated by both Llama-3-8b-Instruct and Llama-3-70b-Instruct using various decoding strategies. We find that there is little difference in lexical diversity (as measured by TTR) and readability (as measured by FK scores) between MBR-generated, BoN-generated and greedy-decoded outputs.
We attempt to identify linguistic differences (beyond generation length) between MBR-generated responses and responses produced using greedy decoding and BoN decoding. We first consider looking for distinctive features in the formatting of MBR-generated responses by identifying responses to AlpacaEval that make use of bullet and numbered lists and then computing the percentage of outputs that fall into each category. We use Llama-3-8b-Instruct and Llama-3-70b-Instruct as the generator models and Prometheus-2-7b as the MBR judge, and document the results in Table 8. We find that there is little difference in the formatting of MBR-generated, BoN-generated and greedy-decoded outputs.
Next, we analyse the lexical diversity and readability of MBR-generated outputs, using the same dataset, generator and judge model as for the formatting experiments. For lexical diversity, we measure the type token ratio (Richards, 1987) of outputs, and for readability we compute Flesch Kincaid readability scores (Flesch, 1948). We document our findings in Table 9, and leave further analysis of linguistic differences between various decoding strategies to future work.
A.6 Speculative Decoding
Avg time per generation (s) Avg tokens / s GPT-4o score
Vanilla MBR 59.3 4.48 6.83
MBR + Spec. Decoding 52.9 5.21 6.90
Table 10: Results demonstrating that MBR decoding can be combined with speculative decoding to increase autoregressive decoding speed. We measure the average time per generation and the average tokens per second associated MBR decoding with and without speculative decoding, and show that speculative decoding improves decoding speed without sacrificing output quality.
We can combine MBR decoding with speculative decoding (Leviathan et al., 2023) to speed up the MBR decoding step. We demonstrate this on 100 samples uniformly drawn from AlpacaEval, using Llama-3-70b-Instruct as both the generator and the MBR judge (see huggingface.co/ibm-fms/Llama-3-70b-accelerator for our chosen draft model), and report our findings in Table 10.
A.7 Rank Correlation with GPT-4o Scores
Avg. $\Delta$ over Greedy Avg. Rank Corr
Prometheus-7b 0.28 0.119
Prometheus-8x7b 0.39 0.136
JudgeLM-7b 0.22 0.053
JudgeLM-33b 0.31 0.113
Llama-3-8b-Instruct 0.28 0.116
Llama-3-70b-Instruct 0.41 0.144
Table 11: Spearman’s rank correlation between the GPT-4o scores and MBR scores generated using various MBR judge models, along with average $\Delta$ over greedy decoding, computed on Llama-3-8b-Instruct outputs on turn 1 of MT-Bench. We find that stronger judges (as measured by the avg. $\Delta$ over greedy) are associated with higher average rank correlations with GPT-4o scores.
We used GPT-4o as a reference-free judge to score the outputs of Llama-3-8b-Instruct ($N_{\mathrm{cand}}=30$, $t=0.7$) on the first turn of MT-Bench. Outputs were scored on a scale of 1 to 10, using the standard MT-bench evaluation approach. We used a GPT-4o judge temperature of 0.5 and generated three scores per generation, taking as our final scores the average of the three sampled scores. Then, for every sample, we computed the Spearman’s rank correlation between the GPT-4o scores and the MBR scores.
We find that stronger judges (higher avg. $\Delta$ over greedy) are generally associated with better correlation with GPT-4o scores, although the absolute value of this correlation is not very high. This suggests that our judges are unlikely to be overly biased towards GPT-4o, although stronger judges will generally agree with GPT-4o more.
@@ Appendix D AlpacaEval Categories
Appendix E Llama-3 as an MBR and BoN Utility Metric
We experiment with using Llama-3-8b-Instruct and Llama-3-70b-Instruct as MBR and BoN utility metrics in Section 3.3.1. This entails prompting the models to act as reference-based or reference-free evaluators. Our prompt for single-turn reference-based evaluation is as follows:
@@ G.3 MBR Self-Training with ROUGE
G.4 MBR Self-Training with Llama-3-8b-Instruct
We also consider using Llama-3-8b-Instruct in place of Prometheus as the utility metric to self-train Llama-3-8b. We compare this approach to using Prometheus to train Llama-3-8b, and find that both approaches lead to significant gains. The fact that Llama-3-8b-Instruct as the judge can be used to improve Llama-3-8b suggests that self-improvement using MBR decoding is possible. We leave this to future work.
Prometheus Llama-3-8b-Instruct
sft 6.70 6.70
dpo-1-MBR 6.94 6.99
dpo-2-MBR 7.45 7.51
dpo-3-MBR 7.55 7.52
Table 16: MT-Bench scores for Llama-3-8b self-trained using DPO with either Prometheus or Llama-3-8b-Instruct as the MBR judge. Generation is done using greedy decoding for all models. We find that both approaches lead to significant gains.
G.5 Alternative Pair Selection Strategies
@@ G.6 Performance by Category
We replicate the question category analysis described in Section 3.3 and Appendix D for our Prometheus MBR self-trained models and report results in Figure 4. We find that performance relative to the sft models improves across almost all question categories, with performance on writing-based categories improving more than on reasoning-based categories. Performance on mathematical reasoning remains unchanged for the 7B model and decreases for the 13B model.
@@ Appendix H Human Study
Table 19: Head-to-head evaluation of Prometheus MBR and Prometheus BoN vs. greedy decoding for Llama-2-70b conducted on the AlpacaEval dataset by human evaluators.
@@
We conducted four head-to-head experiments. The first two experiments, corresponding to experiments in Section 3, were between Prometheus MBR vs. greedy decoding and Prometheus BoN vs. greedy decoding with Llama-2-70b. The next two experiments, corresponding to experiments in Section 4, were between 13B dpo-3-MBR vs. sft and dpo-3-BoN vs. sft. We used 60 samples for each experiment. Volunteers were given an even distribution of samples from each experiment.
@@
Appendix I Algorithms
I.1 MBR Inference
We provide the algorithm for MBR inference in Algorithm 1, complementing the mathematical overview provided in Section 2.
Algorithm 1 MBR Inference
0:  Prompt $x$, generator model $p$, reference-based utility metric $u$, number of candidates $N_{\mathrm{cand}}$, sampling temperature $t$.
0:  MBR output $\hat{y}$
  Initialise $\mathcal{H}_{\mathrm{hyp}}\leftarrow\emptyset$
  for $i\in\{1,2,\dots,N_{\mathrm{cand}}\}$ do
     Sample $y^{(i)}\sim p(\cdot|x)$ with temperature $t$
     Add $y^{(i)}$ to $\mathcal{H}_{\mathrm{hyp}}$ // Form hypothesis set
  end for
  for $y^{(i)}\in\mathcal{H}_{\mathrm{hyp}}$ do
     Compute $\tilde{u}(y^{(i)})=\frac{1}{N_{\mathrm{cand}}}\sum_{j=1}^{N_{\mathrm{cand}}}u(y^{(i)},y^{(j)})$ // Compute expected utility
  end for
  Select $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{H}_{\mathrm{hyp}}}\tilde{u}(y)$
  return  $\hat{y}$
I.2 MBR Distillation
In MBR distillation, we first gather a preference dataset using MBR decoding over a set of input prompts. Given input prompts $X_{k}$ and a base model $\pi_{\theta_{k-1}}$, we first sample $N_{\mathrm{cand}}$ outputs per prompt $y^{(i)}\sim\pi_{\theta_{k-1}}(\cdot|x)$, where $x\in X_{k}$, forming hypothesis set $\mathcal{H}_{\mathrm{hyp}}=\{y^{(1)},y^{(2)},\dots,y^{(N_{\mathrm{cand}})}\}$.
We then compute expected utility for elements in $\mathcal{H}_{\mathrm{hyp}}$
$\tilde{u}(y^{(i)})=\frac{1}{N_{\mathrm{cand}}}\sum_{j=1}^{N_{\mathrm{cand}}}u(y^{(i)},y^{(j)})$ (4)
and form preference pairs using the output that maximises expected utility $\hat{y}^{+}$ and the output that minimises it $\hat{y}^{-}$.
$\hat{y}^{+}=\operatorname*{arg\,max}_{y\in\mathcal{H}_{\mathrm{hyp}}}\tilde{u}(y)$
$\hat{y}^{-}=\operatorname*{arg\,min}_{y\in\mathcal{H}_{\mathrm{hyp}}}\tilde{u}(y)$
We collect these preference pairs and their prompt $(x,\hat{y}^{+},\hat{y}^{-})$ in a dataset we denote $\mathcal{Y}_{k}$.
We then use these preference pairs for DPO (Rafailov et al., 2024) training. In DPO, we minimise the following policy objective:
$\mathcal{L}_{\mathrm{DPO}}=-\mathbb{E}_{(x,\hat{y}^{+},\hat{y}^{-}\sim\mathcal{Y}_{k})}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(\hat{y}^{+}|x)}{\pi_{\mathrm{ref}}(\hat{y}^{+}|x)}-\beta\log\frac{\pi_{\theta}(\hat{y}^{-}|x)}{\pi_{\mathrm{ref}}(\hat{y}^{-}|x)}\right)\right]$ (5)
where $\pi_{\theta}$ is the policy and $\pi_{\mathrm{ref}}$ the reference model. We repeat this process iteratively with $k=1,2,\dots,K$. The initial model $\pi_{\theta_{0}}$ is the base sft model. We choose $K=3$ in our experiments.
We also provide the algorithm for MBR distillation with DPO in Algorithm 2.
Algorithm 2 MBR Distillation with DPO
0:  Prompt sets $X_{1},X_{2},\dots,X_{K}$, sft model $\pi_{\theta_{0}}$, reference-based utility metric $u$, number of candidates $N_{\mathrm{cand}}$, sampling temperature $t$, number of self-training iterations $K$.
0:  Self-trained model $\pi_{\theta_{K}}$
  for $k\in\{1,2,\dots,K\}$ do
     Initialise $\mathcal{Y}_{k}\leftarrow\emptyset$
     for $x\in X_{k}$ do
        Initialise $\mathcal{H}_{\mathrm{hyp}}\leftarrow\emptyset$
        for $i\in\{1,2,\dots,N_{\mathrm{cand}}\}$ do
           Sample $y^{(i)}\sim\pi_{\theta_{k-1}}(\cdot|x)$ with temperature $t$
           Add $y^{(i)}$ to $\mathcal{H}_{\mathrm{hyp}}$ // Form hypothesis set
        end for
        for $y^{(i)}\in\mathcal{H}_{\mathrm{hyp}}$ do
           Compute $\tilde{u}(y^{(i)})=\frac{1}{N_{\mathrm{cand}}}\sum_{j=1}^{N_{\mathrm{cand}}}u(y^{(i)},y^{(j)})$ // Compute expected utility
        end for
        Select $\hat{y}^{+}=\operatorname*{arg\,max}_{y\in\mathcal{H}_{\mathrm{hyp}}}\tilde{u}(y)$ // Select highest scoring output
        Select $\hat{y}^{-}=\operatorname*{arg\,min}_{y\in\mathcal{H}_{\mathrm{hyp}}}\tilde{u}(y)$ // Select lowest scoring output
        Add $(\hat{y}^{+},\hat{y}^{-})$ to $\mathcal{Y}_{k}$ // Form preference pairs
     end for
     Update $\pi_{\theta_{k}}\leftarrow DPO(\pi_{\theta_{k-1}},\mathcal{Y}_{k})$ // DPO training on preference pairs
  end for
  return  $\pi_{\theta_{K}}$
Appendix J Limitations
While our work demonstrates the significant potential of MBR decoding, there are limitations that should be addressed in future research. Firstly, although we demonstrate using existing judge LLMs utility metrics that MBR decoding consistently outperforms BoN decoding, this does not preclude the existence of reference-free metrics that are powerful enough to match or surpass the performance of their direct reference-based counterparts. This relates to a possible broader limitation on the benefits of using consensus quality for output selection, as the consensus solution may not always be the optimal one. We encourage future work to train better utility metrics in order to better understand these limitations. A second limitation of our work is that we do not study the biases introduced by the utility metric. One particularly pernicious form of bias is “reward-hacking” behavior, where the utility metric (likely as a result of its own training) selects outputs that evaluate well on our benchmarks but that are actually worse in quality. While we preclude this from being the case in our experiments via our human study (Appendix H), this does not mean that such pernicious behavior cannot arise in other settings. Finally, we do not study limitations on scalability. Although we show that small judge LLMs (7B) can serve as utility metrics for much larger models (70B), it is likely that weaker utility metrics cease being useful for very strong LLMs and on very complex tasks. Further research is needed to determine when this breakdown occurs.
Appendix K Training and Inference Hyperparameters
@@ I.2 MBR Distillation
Our SFT and DPO hyperparameters for our self-training experiments in Section 4.2 are provided in Tables 21 and 22. We use bf16 mixed precision training with 8xA100 GPUs for all experiments.